OpenStudy (anonymous):

@mathstudent55 Do you think you could help me?

OpenStudy (studygurl14):

What's your question?

OpenStudy (anonymous):

Okay let me upload it okay? Give me 1 minute...

OpenStudy (studygurl14):

ok

OpenStudy (anonymous):

I'm stressing out big time about this.

OpenStudy (studygurl14):

Don't stress. It's fairly simple once you understand it. Hold on a sec, kay?

OpenStudy (studygurl14):

I have to type it out.

OpenStudy (anonymous):

I have like 10 of these problems and I like just want to cry because I don't know how to do any of them.

OpenStudy (anonymous):

Thank you so much. It means a lot :)

OpenStudy (studygurl14):

Okay, so the lateral area of a three-dimensional figure is basically the surface area of the every single surface, EXCEPT the bases. So, int he picture. You see the rectangle part. That's the part you want to find the area of.

OpenStudy (studygurl14):

And we're lucky here. There's a handy-dandy formula for finding the lateral area of a cylinder (which is what this "net" represents)

OpenStudy (anonymous):

Okay.

OpenStudy (studygurl14):

\(\large\tt\ Lateral\ Area\ of\ a\ Cylinder: \) \(\Large\ 2\pi rh\) Where, r = radius of base, and h = height

OpenStudy (anonymous):

so then it would be 2 * 3.14 * 9 * 17 ?

OpenStudy (studygurl14):

They give you the height already, 17 mm. They also give you the diameter of the base, 9 mm. Do you know how to find the radius using the diameter?

OpenStudy (studygurl14):

Not quite @lily1021 9 mm represents the diameter, not the radius.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

oh 9 is the diameter. No I don't know how to find the radius using the diameter.

OpenStudy (studygurl14):

The radius is half the length of the diameter. So what's 9 divided by 2?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

4.5

OpenStudy (studygurl14):

Great. But I just realized something.

OpenStudy (studygurl14):

Since the formula already has a two in it, you don't even have to divide. Because watch:

OpenStudy (studygurl14):

\(\large 2\pi(\frac{9}{2})h\rightarrow 2(\frac{9}{2})\pi(17)\rightarrow 9\pi(17)\rightarrow the~answer\)

OpenStudy (studygurl14):

One last step to the answer. What's 9(17) ?

OpenStudy (studygurl14):

@lily1021

OpenStudy (anonymous):

153?

OpenStudy (anonymous):

or is that wrong?

OpenStudy (studygurl14):

No that's correct. So teh answer is: \(\large 153\pi\)

OpenStudy (studygurl14):

That's all. Pretty simple, right?

OpenStudy (anonymous):

Wow that's actually pretty simple. So because there is already a 2 in the formula I don't have to divide the diameter by 2 anymore?

OpenStudy (studygurl14):

right. Exactly.

OpenStudy (studygurl14):

You did great. :)

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Do you know how I find surface area of a prism? Is there a formula too?

OpenStudy (studygurl14):

The formula for lateral area of cylinder can either be \(\large 2\pi rh\) or \(\large d\pi h\) Depending on the information given

OpenStudy (anonymous):

Okay so if radius isn't given use the bottom and if it is but not the diameter use the top one. Got it.

OpenStudy (studygurl14):

Yes, @lily1021 . There is a formula for surface area of a prism.

OpenStudy (studygurl14):

@lily1021 do you have more questions?

OpenStudy (anonymous):

No but I'm trying to find the surface area formula for my next problem . Don't I add all the faces and then multiply by the height?

OpenStudy (studygurl14):

Um...sorta. How about you close this question and open a new one. Tag me by typing @StudyGurl14

OpenStudy (anonymous):

Okay thank you.

OpenStudy (studygurl14):

anytime :)